# Feasibility of Learning Weighted Automata on a Semiring

### v3

@@ 2.2. Learning with membership and equivalence queries
We are investigating one type of learning done by membership and equivalence queries [4, 8, 7]. In certain circumstances (i.e. by placing conditions on the weighted automaton and the semiring) it is known that one can learn efficiently a weighted automaton computing the function $f$ when provided with finitely many data points $f(w)$ and some steering. Specifically, suppose that there is an oracle who can answer two types of questions about $f$:
@@
A ‘membership’ query is so-named since in the case of the Boolean semiring, this question amounts to asking whether a given word belongs to the rational language recognised by the automaton computing $f$. More generally, a “membership query” (sometimes also called “output query”) provides a mechanism for a learner to gather data points $f(w)$ from which they may attempt to build a hypothesis automaton $\mathcal{H}$ consistent with these data points. An equivalence query provides a mechanism for a learner to test their hypothesis and receive useful feedback, in the form of success or counterexample. Throughout this paper we consider hypothesis automata constructed in a prescribed manner from a partial Hankel matrix, as we shall now explain. What follows has been introduced and studied in the literature; we give proofs of some known results for completeness, putting these into our general framework. Our presentation differs slightly and leads to questions that are answered in Sections 3 and 4 and represent our main contribution.
@@
We will use the notation $\langle x\rangle$ for elements of $S^{\Sigma^{*}}$, and for a word $w\in\Sigma^{*}$ we write $\langle x\rangle_{w}$ to denote the $w$-th component of $\langle x\rangle$. The restriction of $\langle x\rangle$ to a subset $Z\subseteq\Sigma^{*}$ will be denoted by $\langle x\rangle_{Z}$. Given $\langle x\rangle\in S^{\Sigma^{*}}$ and $u\in\Sigma^{*}$, we denote by $\langle xu\rangle$ the element of $S^{\Sigma^{*}}$ defined as the shift of $\langle x\rangle$ by $u$, that is, $\langle xu\rangle_{w}=\langle x\rangle_{uw}$ for all words $w$. Given a function $f:\Sigma^{*}\to S$, the infinite matrix $F$ whose rows and columns are indexed by $\Sigma^{*}$ and with entries $F_{w,w^{\prime}}=f(ww^{\prime})$ is called the Hankel matrix of $f$. If $x$ is a word and $f$ and $F$ are clear from context, we shall write $\langle x\rangle$ to denote the infinite row indexed by $x$ in $F$, viewed as an element of $S^{\Sigma^{*}}$. Note that if $x$ is a word, $f$ a function and $F$ its Hankel matrix, this definition is consistent: $\langle xu\rangle_{w}=F_{xu,w}=f(xuw)=F_{x,uw}=\langle x\rangle_{uw}$. For a subset $X\subseteq S^{\Sigma^{*}}$ and $\langle x\rangle$ in $X$, we will sometimes simplify notation by writing $x\in X$.
@@
Let $Z\subseteq\Sigma^{*}$ and $f:\Sigma^{*}\rightarrow S$. A subset $X\subseteq S^{\Sigma^{*}}$ is said to be: (i) left-closed if for all $\langle x\rangle$ in $X$ and $a\in\Sigma$, $\langle xa\rangle$ is left-generated by $X$; (ii) left-closed on $Z$ if for all $\langle x\rangle$ in $X$ and $a\in\Sigma$, $\langle xa\rangle_{Z}$ is left-generated by the elements of $X$ restricted to $Z$; (iii) row-generating for $f$ if every row of the Hankel matrix of $f$ is left-generated by $X$. By an abuse of notation, we say that a set of words [^1] Using $Q$ as the label for a set of words might seem odd at this point, but the reason behind this will become apparent very shortly. $Q$ is left-closed (resp. left-closed on $Z$, row-generating for $f$) if the set $X=\{\langle q\rangle:q\in Q\}$ has this property.
Observe that for any subset $X\subseteq S^{\Sigma^{*}}$, if $X$ is left-closed and left-generates the row $\langle\varepsilon\rangle$ of the Hankel matrix of $f$ then $X$ is row-generating for $f$ (indeed: since $X$ left generates $\langle\varepsilon\rangle$ we may write $\langle\varepsilon\rangle=\bigoplus_{\langle x\rangle\in X}\lambda_{\langle x\rangle}\otimes\langle x\rangle$, with only finitely many coefficients non-zero. Then for all words $u,v\in\Sigma^{*}$ we have $\langle u\rangle_{v}=\langle\varepsilon\rangle_{uv}=\bigoplus_{\langle x\rangle\in X}\lambda_{\langle x\rangle}\otimes\langle x\rangle_{uv}$, hence giving $\langle u\rangle_{v}=\bigoplus_{\langle x\rangle\in X}\lambda_{\langle x\rangle}\otimes\langle xu\rangle_{v}$. Using the fact that $X$ is left-closed, an easy induction on the length of $u$ then gives that $\langle xu\rangle$ is left-generated by $X$, showing that each row $\langle u\rangle$ is a left-linear combination of elements of $X$.) Conversely, if $X\subseteq S^{\Sigma^{*}}$ is a row-generating set for $f$, then we clearly have in particular that $X$ left-generates the row $\langle\varepsilon\rangle$, and in the case where $X$ is a subset of the rows of the Hankel matrix of $f$, we also have that $X$ is left-closed (since we may view the elements $\langle x\rangle\in X$ in correspondence with words $x\in\Sigma^{*}$, in which case $xa\in\Sigma^{*}$ and hence $\langle xa\rangle$ is left-generated by $X$).
@@
It turns out that if one can find a finite subset $X\subseteq S^{\Sigma^{*}}$ that is left-closed and such that the row $\langle\varepsilon\rangle$ is left-generated by $X$, then one can construct an automaton computing the correct function. We call this a Hankel automaton and describe its construction below. However, in practice, using finitely many membership queries only a finite submatrix of the Hankel matrix can be obtained. A hypothesis automaton is defined in the same way as a Hankel automaton, replacing the finite subset $X$ by a finite set of rows of the Hankel matrix (i.e. indexed by a set of words) restricted to a finite set of columns. For Boolean automata, it is always possible to find a finite set amongst the rows of the Hankel matrix and a finite set of columns such that the corresponding hypothesis automaton computes the target function. The $L^{*}$-algorithm is based on this idea.
@@
Notation 2.5.
Given a function $f:\Sigma^{*}\rightarrow S$, a finite subset $Q$ of $S^{\Sigma^{*}}$ and a subset $T$ of $\Sigma^{*}$, we define $\Lambda_{Q,T}\subseteq S^{Q}\times S^{Q\times\Sigma\times Q}$ to be the set of tuples $\lambda=(\lambda_{q}^{\varepsilon},\lambda_{q,a,p})_{q\in Q,a\in\Sigma,p\in Q}$ such that:
  • $\langle\varepsilon\rangle_{T}=\bigoplus_{q\in Q}\lambda_{q}^{\varepsilon}\otimes\langle q\rangle_{T}$, and
  • for all $q\in Q$ and all $a\in\Sigma$, $\langle qa\rangle_{T}=\bigoplus_{p\in Q}\lambda_{q,a,p}\otimes\langle p\rangle_{T}$.
@@
Note that the set $\Lambda_{Q,T}$ depends on the function $f$ (and hence also the semiring $S$), although we have chosen to suppress this dependence in our notation for readability. An element $\lambda\in\Lambda_{Q,T}$ represents a collection of coefficients witnessing that $\langle\varepsilon\rangle_{T}$ and each of the $\langle qa\rangle_{T}$ (for all $q\in Q$ and all $a\in\Sigma$) is a left-linear combinations of the $\langle q\rangle_{T}$ for $q$ in $Q$. In other words, $\Lambda_{Q,T}$ represents the solution space to the system of $S$-linear equations outlined in Notation 2.5. The set $\Lambda_{Q,T}$ is non-empty if and only if a solution to this system exists; in the language introduced above, if and only if $Q$ is left-closed on $T$ and $\langle\varepsilon\rangle_{T}$ is left-generated by $Q$ restricted to $T$. Note also that if $T\subseteq T’$ ’, then $\Lambda_{Q,T}\supseteq\Lambda_{Q,T’}$. If $T=\Sigma^{*}$, we simply write $\Lambda_{Q}$.
@@
Given a finite subset $Q$ of $S^{\Sigma^{*}}$ and a subset $T$ of $\Sigma^{*}$, for each $\lambda\in\Lambda_{Q,T}$, we define a finite weighted automaton $\mathcal{H}_{Q,T,\lambda}$ with:
@@
  • initial-state vector $(\lambda^{\varepsilon}_{q})_{q\in Q}$;
  • final-state vector $(\langle q\rangle_{\varepsilon})_{q\in Q}$; and
  • for all $q,p\in Q$ and all $a\in\Sigma$, a transition from $q$ to $p$ labelled by $a$ with weight $\lambda_{q,a,p}$.
If $T=\Sigma^{*}$, we say $\mathcal{H}_{Q,T,\lambda}$ is a Hankel automaton and write simply $\mathcal{H}_{Q,\lambda}$ to reduce notation. If $Q$ is a finite set of rows of the Hankel matrix and $T$ is finite, we say that $\mathcal{H}_{Q,T,\lambda}$ is a hypothesis automaton, and view $Q$ as a finite set of words by identifying with an appropriate index set.
@@
Consider the function $f$ over $\mathbb{N}_{\max}$ computed by our running example from Figure 1(a). The set $\{\langle\varepsilon\rangle,\langle a\rangle,\langle b\rangle\}$ clearly left-generates $\langle\varepsilon\rangle$ and, as shown previously, is left-closed. Let $q_{\varepsilon},q_{a}$ and $q_{b}$ denote these rows. Then considering the coefficients $\lambda^{\varepsilon}_{q_{\varepsilon}}=0$, $\lambda_{q_{\varepsilon},a,q_{a}}=\lambda_{q_{\varepsilon},b,q_{b}}=\lambda_{q_{a},b,q_{a}}=\lambda_{q_{b},a,q_{b}}=0$, $\lambda_{q_{a},a,q_{a}}=\lambda_{q_{b},b,q_{b}}=1$ and all non-specified coefficients $\infty$- ∞, the corresponding Hankel automaton would be obtained from the one in Figure 1(a) by replacing the weight of the transitions starting at $q_{\varepsilon}$ by 00, and replacing the final weights by $1$. For $Q=\{q_{\varepsilon},q_{a}\}$ and $T=\{\varepsilon\}$ we have also seen that $Q$ is left-closed on $T$. By taking $\lambda^{\varepsilon}_{q_{\varepsilon}}=0$, $\lambda_{q_{\varepsilon},a,q_{a}}=\lambda_{q_{\varepsilon},b,q_{a}}=\lambda_{q_{a},b,q_{a}}=0$, $\lambda_{q_{a},a,q_{a}}=1$ and all remaining coefficients $\infty$- ∞, one obtains a hypothesis automaton with two states. Notice that this automaton does not compute $f$; it computes $|w|_{a}$ if $w$ starts with an $a$, $|w|_{a}+1$ if $w$ starts with a $b$, and $\infty$- ∞ if $w=\varepsilon$. Note that if $\{\varepsilon,aw\}\subseteq T^{\prime}$ for some word $w$, then $Q=\{q_{\varepsilon},q_{a}\}$ is not left-closed on $T^{\prime}$.
In the case where $S$ is a field the following result can be found in [19]. However, the proof can easily be adapted to the case of arbitrary semirings, and since the details are instructive, we include a short proof.
@@
Let $f:\Sigma^{*}\rightarrow S$ be a function. There exists a finite subset $Q\subseteq S^{\Sigma^{*}}$ that is left-closed and left-generates $\langle\varepsilon\rangle$ if and only if $f$ is computed by a finite-state weighted automaton $\mathcal{A}$ over $S$. Specifically, if $Q\subseteq S^{\Sigma^{*}}$ is a finite left-closed set that left-generates $\langle\varepsilon\rangle$, then $\mathcal{H}_{Q,\lambda}$ computes $f$ for all $\lambda\in\Lambda_{Q}\neq\emptyset$, whilst if $f$ is computed by the finite-state weighted automaton $\mathcal{A}$ over $S$, then the set of all $\langle q\rangle$ (where $q$ is a state of $\mathcal{A}$ and $\langle q\rangle_{w}$ is defined as the value of $w$ in $\mathcal{A}$ if $q$ was considered the unique initial state with weight $1_{S}$ and all other weights are the same) is left-closed and left-generates every row of the Hankel matrix of $f$.
@@
Before presenting the proof, we alert the reader to the fact that in Theorem 2.9 the set $Q$ is not assumed to be a set of rows of the Hankel matrix of $f$. In light of Remark 2.6, if $Q$ is a finite set of rows of the Hankel matrix of $f$ and $\Lambda_{Q}$ is non-empty, then for any $\lambda\in\Lambda_{Q}$ the Hankel automaton $\mathcal{H}_{Q,\lambda}$ computes the function $f$. For any finite set $T\subseteq\Sigma^{*}$ we then necessarily have $\lambda\in\Lambda_{Q,T}$, and so (by construction) the hypothesis automaton $\mathcal{H}_{Q,T,\lambda}$ constructed from the finite $(Q,T)$ submatrix of the Hankel matrix of $f$, will also compute $f$. The success of Angluin-type algorithms in the existing literature are based on the ability to find (using finitely many membership and equivalence queries): a finite set of words $Q$ such that $\Lambda_{Q}$ is non-empty, together with a finite set of words $T$ such that $\Lambda_{Q,T}=\Lambda_{Q}$.

### v4

@@ 2.2. Learning with membership and equivalence queries
We are investigating one type of learning done by membership and equivalence queries [4, 9, 7] where the learner makes guesses by solving systems of left-linear (or right-linear) equations over $S$. Specifically, suppose that there is an oracle who can answer two types of questions about a function $f:\Sigma^{*}\rightarrow S$:
@@
A ‘membership’ query is so-named since in the case of the Boolean semiring, this question amounts to asking whether a given word belongs to the rational language recognised by the automaton computing $f$. More generally, a “membership query” (sometimes also called “output query”) provides a mechanism for a learner to gather data points $f(w)$ from which they may attempt to build an hypothesis automaton $\mathcal{H}$ consistent with these data points. An equivalence query provides a mechanism for a learner to test their hypothesis and receive useful feedback, in the form of success or counterexample. Throughout this paper we consider hypothesis automata constructed in a prescribed manner from a partial Hankel matrix, as we shall explain below. In certain circumstances (i.e. by placing conditions on the weighted automaton and/or the semiring) it is already known that one can efficiently learn a weighted automaton computing the function $f$ using finitely membership and equivalence queries. Indeed, the underlying ideas introduced in the remainder of this section have been introduced and studied in the literature; we give proofs of some known results for completeness, putting these into our general framework. Our presentation differs slightly and leads to questions that are answered in Sections 3 and 4 and represent our main contribution.
@@
As before, we use the notation $\underline{x}$ for elements of $S^{\Sigma^{*}}$, and for a word $w\in\Sigma^{*}$ we write $\underline{x}_{w}$ to denote the $w$-th component of $\underline{x}$. The restriction of $\underline{x}$ to a subset $Z\subseteq\Sigma^{*}$ will be denoted by $\underline{x}_{Z}$. We shall also make use of the following right action of $\Sigma^{*}$ on $S^{\Sigma^{*}}$: given $\underline{x}\in S^{\Sigma^{*}}$ and $u\in\Sigma^{*}$, we denote by $\underline{x}\cdot u$ the element of $S^{\Sigma^{*}}$ defined as the shift of $\underline{x}$ by $u$, that is, $(\underline{x}\cdot u)_{w}=\underline{x}_{uw}$ for all words $w$. The right action of $\Sigma$ on $S^{\Sigma^{*}}$ commutes with the left action of $S$ on $S^{\Sigma^{*}}$ in the sense that for all $\lambda\in S$, $\underline{x}\in S^{\Sigma^{*}}$ and $u\in\Sigma^{*}$ we have $(\lambda\cdot\underline{x})\cdot u=\lambda\cdot(\underline{x}\cdot u)$. Given a function $f:\Sigma^{*}\to S$, the infinite matrix $F$ whose rows and columns are indexed by $\Sigma^{*}$ and with entries $F_{w,w^{\prime}}=f(ww^{\prime})$ is called the Hankel matrix of $f$. It will also be convenient to have a notation for elements of $S^{\Sigma^{*}}$ that arise as rows of the Hankel matrix. If $v\in\Sigma^{*}$ is a word and $f$ and $F$ are clear from context, we shall write $\langle v\rangle$ to denote the infinite row indexed by $v$ in $F$, viewed as an element of $S^{\Sigma^{*}}$. Note that if $v$ is a word, $f$ a function and $F$ its Hankel matrix, then the right action of $\Sigma^{*}$ on $\langle v\rangle$ corresponds to the obvious right action of $\Sigma^{*}$ on $v$, that is: $(\langle v\rangle\cdot u)_{w}=(\langle v\rangle)_{uw}=F_{v,uw}=f(vuw)=F_{vu,w}=\langle vu\rangle_{w}$.
@@
Let $Z\subseteq\Sigma^{*}$ and $f:\Sigma^{*}\rightarrow S$. A subset $X\subseteq S^{\Sigma^{*}}$ is said to be: (i) left-closed if for all $\underline{x}$ in $X$ and $a\in\Sigma$, $\underline{x}\cdot a$ is left-generated by $X$; (ii) left-closed on $Z$ if for all $\underline{x}$ in $X$ and $a\in\Sigma$, $(\underline{x}\cdot a)_{Z}$ is left-generated by the elements of $X$ restricted to $Z$; (iii) row-generating for $f$ if every row of the Hankel matrix of $f$ is left-generated by $X$. By an abuse of language, we say that a set of words [^1] Using $Q$ as the label for a set of words might seem odd at this point, but the reason behind this will become apparent very shortly. $Q$ is left-closed (resp. left-closed on $Z$, row-generating for $f$) if the set $\underline{Q}=\{\langle q\rangle:q\in Q\}$ has this property.
The following simple observations are key in what follows.
Lemma 2.4.
Let $X\subseteq S^{\Sigma^{*}}$.
  1. (1)
    If $X$ is left-closed and left-generates the row $\langle\varepsilon\rangle$ of the Hankel matrix of $f$ then $X$ is row-generating for $f$.
  2. (2)
    If $X$ is a subset of the rows of the Hankel matrix and is a row-generating set for $f$, then $X$ is left-closed (and left-generates the row $\langle\varepsilon\rangle$).
Proof.
(1) Since $X$ left generates $\langle\varepsilon\rangle$ we may write $\langle\varepsilon\rangle=\bigoplus_{\underline{x}\in X}\lambda_{\underline{x}}\otimes\underline{x}$, where only finitely many of the coefficients $\lambda_{\underline{x}}$ are non-zero. Then for all words $u,v\in\Sigma^{*}$ we have
$\langle u\rangle_{v}=\langle\varepsilon\rangle_{uv}=\bigoplus_{\underline{x}\in X}\lambda_{\underline{x}}\otimes\underline{x}_{uv}=\bigoplus_{\underline{x}\in X}\lambda_{\underline{x}}\otimes(\underline{x}\cdot u)_{v}.$
Then, using the fact that $X$ is left-closed, an easy induction on the length of $u$ then gives that $\underline{x}\cdot u$ is left-generated by $X$, in turn showing that each row $\langle u\rangle$ of the Hankel matrix is a left-linear combination of elements of $X$.
 (2) That $X$ left-generates $\langle\varepsilon\rangle$ is completely clear from definition of row generating set. Since $X$ is a subset of the rows of the Hankel matrix of $f$, each element of $X$ may be written in the form $\langle w\rangle$ for some $w\in\Sigma^{*}$, in which case $wa\in\Sigma^{*}$ for all $a\in\Sigma$, and hence $\langle w\rangle\cdot a=\langle wa\rangle$ is left-generated by $X$. ∎
@@
It turns out that if one can find a finite subset $X\subseteq S^{\Sigma^{*}}$ that is left-closed and such that the row $\langle\varepsilon\rangle$ is left-generated by $X$, then by solving the corresponding system of $S$-linear equations involving the (infinite) vectors in $X\cup\{\langle\varepsilon\rangle\}$ one can construct an automaton computing the correct function. We call this a Hankel automaton and describe its construction below. However, in practice, using finitely many membership queries only a finite submatrix of the Hankel matrix can be obtained. An hypothesis automaton is defined in the same way as a Hankel automaton, replacing the finite subset $X$ by a finite set of rows of the Hankel matrix (i.e. indexed by a set of words) restricted to a finite set of columns. For Boolean automata, it is always possible to find a finite set amongst the rows of the Hankel matrix and a finite set of columns such that the corresponding hypothesis automaton computes the target function. The $L^{*}$-algorithm is based on this idea.
@@
Notation 2.6.
Given a function $f$ (and in particular the row indexed by $\varepsilon$ in its Hankel matrix), a finite non-empty subset $Q$ of vectors of $S^{\Sigma^{*}}$ and a non-empty subset $T$ of $\Sigma^{*}$, we define the left solution set $\Lambda_{Q,T}$ to be the set of vectors $\underline{\lambda}\in S^{Q\cup(Q\times\Sigma\times Q)}$ satisfying:
  • $\langle\varepsilon\rangle_{T}=\bigoplus_{\underline{q}\in Q}\underline{\lambda}_{\underline{q}}\otimes\underline{q}_{T}$, and
  • for all $\underline{q}\in Q$ and all $a\in\Sigma$, $(\underline{q}\cdot a)_{T}=\bigoplus_{\underline{p}\in Q}\underline{\lambda}_{\underline{q},a,\underline{p}}\otimes\underline{p}_{T}$.
If $T=\Sigma^{*}$, we will simply write $\Lambda_{Q}$ instead of $\Lambda_{Q,\Sigma^{*}}$. Note that the left solution sets $\Lambda_{Q,T}$ and $\Lambda_{Q}$ depend on the function $f$ (and hence also the semiring $S$), although we have chosen to suppress this dependence in our notation for readability. If $W$ is a finite subset of $\Sigma^{*}$ we will also, by a slight abuse of notation, write $\Lambda_{W}$ and $\Lambda_{W,T}$ to mean $\Lambda_{Q}$ and $\Lambda_{Q,T}$ where $Q=\{\langle w\rangle:w\in W\}$.
@@
An element $\underline{\lambda}\in\Lambda_{Q,T}$ represents a collection of coefficients witnessing that $\langle\varepsilon\rangle_{T}$ and each of the $(\underline{q}\cdot a)_{T}$ (for all $\underline{q}\in Q$ and all $a\in\Sigma$) is a left-linear combinations of the $\underline{q}_{T}$ for $q$ in $Q$. In other words, $\Lambda_{Q,T}$ represents the solution space to the system of $S$-linear equations outlined in Notation 2.6. The key idea here is that the set $\Lambda_{Q,T}$ is non-empty if and only if a solution to this system exists; in the language introduced above, if and only if $Q$ is left-closed on $T$ and $\langle\varepsilon\rangle_{T}$ is left-generated by $Q$ restricted to $T$.
Remark 2.8.
Let $T\subseteq\Sigma^{*}$, and let $Q\subseteq S^{\Sigma^{*}}$ be a finite set. If $T\subseteq T’\subseteq\Sigma^{*}$, then it is clear from the above equations that $\Lambda_{Q,T}\supseteq\Lambda_{Q,T’}\supseteq\Lambda_{Q}$. However, if $Q^{\prime}$ is a finite set contained in or contained by $Q$, then there is no containment between $\Lambda_{Q,T}$ (which lies in $S^{Q\cup(Q\times\Sigma\times Q)}$) and $\Lambda_{Q^{\prime},T}$ (which lies in $S^{Q^{\prime}\cup(Q^{\prime}\times\Sigma\times Q^{\prime})}$).
Lemma 2.9.
Let $f:\Sigma^{*}\rightarrow S$ be a function, $T\subseteq\Sigma^{*}$, and suppose that $Q^{\prime}\subseteq Q$ are finite sets of rows of the Hankel matrix of $f$. If $\Lambda_{Q^{\prime}}\neq\emptyset$ then $\Lambda_{Q,T}\neq\emptyset$.
Proof.
Since each element of $Q$ is a row of the Hankel matrix, there exists a finite set of words $W\subseteq\Sigma^{*}$ such that $Q=\{\langle w\rangle:w\in W\}$. Since $\Lambda_{Q^{\prime}}\neq\emptyset$, we have that $Q^{\prime}$ is left-closed and left-generates the row $\langle\varepsilon\rangle$, and so $Q^{\prime}$ is row-generating for $f$, by Lemma 2.4. In particular, $\langle\varepsilon\rangle$ and each row $\langle wa\rangle$ where $w\in W$ and $a\in\Sigma$ can be expressed as a left linear combination of the elements of $Q^{\prime}$, and hence also of $Q$. Since $\langle wa\rangle=\langle w\rangle\cdot a$, it then follows easily that the system of equations given in Notation 2.6 has a solution. ∎
@@
Let $f:\Sigma^{*}\rightarrow S$ be a function, and suppose that $Q$ is a finite subset of $S^{\Sigma^{*}}$ and $T$ a subset of $\Sigma^{*}$ such that $\Lambda_{Q,T}\neq\emptyset$. Then for each $\underline{\lambda}\in\Lambda_{Q,T}$, we define a finite weighted automaton $\mathcal{H}_{Q,T,\underline{\lambda}}$ with:
@@
  • initial-state vector $\underline{\lambda}_{Q}$;
  • final-state vector $(\underline{q}_{\varepsilon})_{\underline{q}\in Q}$; and
  • for all $\underline{q},\underline{p}\in Q$ and all $a\in\Sigma$, a transition from state $\underline{q}$ to state $\underline{p}$ labelled by $a$ with weight $\underline{\lambda}_{\underline{q},a,\underline{p}}$.
If $Q$ is a finite set of rows of the Hankel matrix of $f$ and $T=\Sigma^{*}$, we say $\mathcal{H}_{Q,T,\underline{\lambda}}$ is a Hankel automaton and write simply $\mathcal{H}_{Q,\underline{\lambda}}$ to reduce notation. If $Q$ is a finite set of rows of the Hankel matrix and $T$ is finite, we say that $\mathcal{H}_{Q,T,\underline{\lambda}}$ is an hypothesis automaton; in this case we may view $Q$ as a finite set of words by identifying with an appropriate index set (note: there need not be a unique set of words).
@@
Consider the function $f$ over $\mathbb{N}_{\max}$ computed by our running example from Figure 1(a). The set $\{\langle\varepsilon\rangle,\langle a\rangle,\langle b\rangle\}$ containing three rows of the Hankel matrix clearly left-generates $\langle\varepsilon\rangle$ and, as shown previously, is left-closed. Let $q_{\varepsilon},q_{a}$ and $q_{b}$ denote these rows (omitting the underlining of these elements for ease of reading). Then considering the coefficients $\underline{\lambda}_{q_{\varepsilon}}=0$, $\underline{\lambda}_{q_{\varepsilon},a,q_{a}}=\underline{\lambda}_{q_{\varepsilon},b,q_{b}}=\underline{\lambda}_{q_{a},b,q_{a}}=\underline{\lambda}_{q_{b},a,q_{b}}=0$, $\underline{\lambda}_{q_{a},a,q_{a}}=\underline{\lambda}_{q_{b},b,q_{b}}=1$ and all non-specified coefficients $\infty$- ∞, the corresponding Hankel automaton would be the automaton with three states obtained from the one in Figure 1(a) by replacing the weight of the transitions starting at $q_{\varepsilon}$ by 00, and replacing the final weights by $1$. For $Q=\{q_{\varepsilon},q_{a}\}$ and $T=\{\varepsilon\}$ we have also seen that $Q$ is left-closed on $T$. By taking $\underline{\lambda}_{q_{\varepsilon}}=0$, $\underline{\lambda}_{q_{\varepsilon},a,q_{a}}=\underline{\lambda}_{q_{\varepsilon},b,q_{a}}=\underline{\lambda}_{q_{a},b,q_{a}}=0$, $\underline{\lambda}_{q_{a},a,q_{a}}=1$ and all remaining coefficients $\infty$- ∞, one obtains an hypothesis automaton with two states. Notice that this automaton does not compute $f$; it computes $|w|_{a}$ if $w$ starts with an $a$, $|w|_{a}+1$ if $w$ starts with a $b$, and $\infty$- ∞ if $w=\varepsilon$. Note that if $\{\varepsilon,aw\}\subseteq T^{\prime}$ for some word $w$, then $Q=\{q_{\varepsilon},q_{a}\}$ is not left-closed on $T^{\prime}$.
Fliess’s theorem [21] states that a function is computed by a weighted automaton over the field of reals if and only if its Hankel matrix has finite rank, and moreover, this rank is exactly the minimal number of states of a weighted automaton computing the function. There are many equivalent definitions of the rank of a matrix with entries over a field whose analogues cease to coincide over an arbitrary semirings, however, the proof of Fliess can be easily adapted to obtain the following statement, and since the details are instructive, we include a short proof.
@@
Let $f:\Sigma^{*}\rightarrow S$ be a function. The following are equivalent:
  1. (1)
    there exists a finite subset $Q\subseteq S^{\Sigma^{*}}$ that is both left-closed and left-generates the row indexed by $\varepsilon$ in the Hankel matrix of $f$ (i.e. $\Lambda_{Q}\neq\emptyset$ for some finite subset $Q\subseteq S^{\Sigma^{*}}$);
  2. (2)
    $f$ is computed by a finite-state weighted automaton $\mathcal{A}$ over $S$.
Specifically, if $\Lambda_{Q}\neq\emptyset$, then $\mathcal{H}_{Q,\underline{\lambda}}$ computes $f$ for all $\underline{\lambda}\in\Lambda_{Q}\neq\emptyset$, whilst if $f$ is computed by the finite-state weighted automaton $\mathcal{A}$ over $S$ with state set $Q$, then $\Lambda_{\underline{Q}}\neq\emptyset$ for $\underline{Q}=\{\underline{q}:q\in Q\}\subseteq S^{\Sigma^{*}}$ where for each $w\in\Sigma^{*}$ we define $\underline{q}_{w}$ to be the value on input $w$ computed by the automaton $\mathcal{A}_{q}$ obtained from $\mathcal{A}$ by making $q$ the unique initial state with weight $1_{S}$ and all other weights are the same.
@@
Before presenting the proof, we alert the reader to the fact that in Theorem 2.12 the set $Q$ is not assumed to be a set of rows of the Hankel matrix of $f$, and as we shall see below, in general, a function computed by a finite-state weighted automaton over an arbitrary semiring may or may not satisfy $\Lambda_{Q}\neq\emptyset$ for some finite subset $Q$ of the rows of its Hankel matrix. However, in view of Notation 2.6, Remark 2.7 and Definition 2.10, if there exists a finite set $Q$ of rows of the Hankel matrix such that $\Lambda_{Q}$ is non-empty, then for any finite set $T\subseteq\Sigma^{*}$ and any $\underline{\lambda}\in\Lambda_{Q}$ we have $\underline{\lambda}\in\Lambda_{Q}\subseteq\Lambda_{Q,T}$, so that (by construction) the hypothesis automaton $\mathcal{H}_{Q,T,\underline{\lambda}}$ constructed from the finite $(Q,T)$ submatrix of the Hankel matrix of $f$, will be equal to the Hankel automaton $\mathcal{H}_{Q,\underline{\lambda}}$ and, so by Theorem 2.12, will compute $f$. The success of Angluin-type algorithms in the existing literature is therefore typically based on the ability to find (using finitely many membership and equivalence queries): a finite set of words $Q$ such that $\Lambda_{Q}$ is non-empty, together with a finite set of words $T$ such that $\Lambda_{Q,T}=\Lambda_{Q}$ in the given setting.